Calculate the NBOD, CBOD, and ThOD of a waste cont
Answer #1

Nitrogenous Biochemical Oxygen Demand is the amount of oxygen need to breakdown this organic nitrogen compound to NO3

NH3+O2----> NO3--+H+ H2O

100 mg/L =100*10-3 g/L

4.57 mgO2 are required for each mgNH3-N oxidized. Hence

for 100mgL =(4.57mgO2/mgNH3-N)*100= 457 mgO2/L

ThOD= 114 mg/L

C3H7OH+O2----> CO2+ H2O

balanced reaction is C3H7OH+ 9/2O2----> 3CO2+4H2O

molecular wegiht of C3H7OH= 36+8+16=60 oxygen required =4.5 moles

ThCOD=4.5 ( 100mg/l/60)=7.5 mg/l

CH3NH2 + 9 /4O2 ----> CO2 + 10/4 H2O + 2/4 N2

Total oxygen COD=

mass of 1mole of CH3NH2= 12+5+14=31

oxygen required =9/4v moles =2.25moles= 2.25*32=72 gms

ThCOD=(100mg/L/31)*72=232